Solve by the elimination method.
.05x+.25y=44
.15x+.05y=48

Answer by checkley77(12844)

.05x+.25y=44 multiply by -3 & add
.15x+.05y=48
-.15x-.75y=-132
--------------------
-.70y=-84
y=-84/-.70
y=120 ans.
,05x+.25*120=44
.05x+30=44
.05x=44-30
.05x=14
x=14/.05
x=280 ans.
Proof:
.15*280+.05*120=48
42+6=48
48=48